# Prediction Sets and Conformal Inference with

### v3

@@ 2 Setup and definitions
A prediction set is defined as a set $C\subset\mathcal{X}\times\mathcal{Y}$, with the interpretation that, given a realisation of the predictor $X=x$, we predict that $Y$ is likely to fall inside the section $C(x)=\left\{y:(x,y)\in C\right\}$. In the absence of interval censoring, estimation of a prediction set that satisfies certain optimality criteria (i.e. the minimal prediction set with coverage guarantees) has been studied in the literature; see Lei and Wasserman, (2014), Barber et al., (2021) among others. As we discuss later in this section, the oracle prediction set without interval censoring is related to the upper level set of the density function of $Y$ given $X$ under point identification.
The situation is significantly different when we are faced with interval censoring and $P_{X,Y}$ is partially identified. In the rest of this section, we first define an optimality criterion for a prediction set under censoring, and analyse why the classical oracle prediction set (which corresponds to an upper level set) is no longer applicable under censoring. We then propose a feasible estimation procedure for the oracle prediction set. We discuss the potential reasons for why the procedure can result in conservative prediction sets and show that the proposed procedure does not incur any unnecessary conservativeness.
There are two criteria for an oracle prediction set: validity and efficiency (see Vovk et al.,, 2009; Lei et al.,, 2013). Let $\mathcal{P}_{I}$ be the identified set of the joint distribution of $(X,Y)$ under censoring. That is, $\mathcal{P}_{I}$ is the set of joint distributions of $(X,Y)$ that are compatible with the joint distribution $P_{X,Y^{L},Y^{U}}$ and some censoring mechanism. A prediction set $C$ is valid if it satisfies the following coverage properties for a new sample $(X_{n+1},Y_{n+1})$ drawn from the same distribution $P$ at a given miscoverage level $\alpha$.
@@
When the model is point identified, $\mathcal{P}_{I}=\left\{P_{X,Y}\right\}$ is a singleton set, the conditions in Equations (2) and (3) reduce to the standard conditions of validity, see for example, Lei and Wasserman, (2014). We will simply refer to these two conditions as marginal validity and conditional validity henceforth.
It is natural to require a prediction set to be valid, so that we have confidence it will contain the true value of the outcome with a certain probability. However, the requirement of validity by itself is vacuous. For example, the prediction set $C$ such that $C(x)=(-\infty,\infty)$, for all $x\in\mathcal{X}$ is technically valid at any miscoverage level but it is also uninformative. Therefore, our target parameter of interest is an oracle prediction set, denoted by $C^{*}_{\mathcal{P}_{I}}$ which is defined as the minimal-volume prediction set that satisfies the validity conditions in Definition 2.1. For simplicity, we will omit the subscript and write $C^{*}=C^{*}_{\mathcal{P}_{I}}$ whenever the dependence on the identified set ${\mathcal{P}_{I}}$ is clear. Since the minimal-volume prediction set is intuitively the most informative among all valid prediction sets, this property is also called efficiency. [^1] This notion of efficiency, (see Vovk et al.,, 2009; Lei et al.,, 2013; Lei and Wasserman,, 2014) differs from the usual definition, where an estimator is considered efficient if it has the smallest variance within a class of estimators. However, the concepts are related, as an estimator with smaller variance generally leads to the construction of tighter and more informative confidence intervals.
@@
$\mathcal{C}=\left\{\bigsqcup_{m=1}^{M}[a_{m},b_{m}]:a_{m}\leq b_{m}<a_{m+1},M<\infty\right\}$
denotes the collection of all finite disjoint union of closed intervals, and the symbol $\sqcup$ indicates that the union is disjoint. The following oracle prediction set under partial identification will be our estimation target.
@@
The oracle prediction set under point identification is $C^{*}_{P_{Y|X}}=\left\{(x,y):x\in\mathcal{X},y\in L(x,\lambda_{x}^{\alpha})\right\}$, which is both conditionally and marginally valid (see Lei and Wasserman, (2014)). For a given $x$, it collects the values of $y$ with sufficiently high density and intuitively assembles the set of $y$’s with the highest probability for a given volume. Once this is calibrated to yield the desired miscoverage level, one obtains a minimal-volume set. Estimating the oracle prediction set $C^{*}_{P_{Y|X}}$ therefore reduces to the problem of estimating the upper level sets $L(x,\lambda)$ of the conditional density $p(y\mid x)$ and selecting an appropriate threshold parameter.
However, this characterisation of the oracle prediction set in terms of level sets is no longer applicable under interval censoring as $P_{Y\mid X}$ is not identified. In fact, the next proposition states that no meaningful bounds can be obtained for the level set of the conditional density of $Y$ given $X$ under interval censoring. For brevity, we temporarily drop the conditioning predictors $X$ in the following proposition. (It can be interpreted as conditional on $X=x$ for $x\in\mathcal{X}$.)
@@
Lemma 2.2 (Theorem SIR-2.3 in Molinari, (2020)).
@@
for a disjoint union of intervals $C\in\mathcal{C}$. As a result, a prediction set that satisfies the feasible version of the constraint will have coverage guarantee since
@@
The above lemma shows that even though the oracle prediction set can be conservative, i.e. overcovering $Y$ for the distribution actually generating the data, the conservativeness is only due to partial identification and the conditional distribution of $(Y^{L},Y^{U})$ and not because of the constraint we focus on differs from the original one.
@@
The oracle prediction set $C^{*}$ is the solution to the following feasible optimisation problem
@@
A natural estimator $\hat{C}$ for the oracle prediction set can be obtained by solving the sample version of the optimisation problem above. The next two sections study the consistency of such estimators for the oracle prediction set, and provides a modification of the estimated prediction set that has finite sample coverage guarantees using conformal inference.
@@ 3 Consistency
We first consider the estimation of the oracle prediction interval $C^{*}_{I}=[\tau_{0}(x),\tau_{1}(x)]$, for $x\in\mathcal{X}$. This analysis highlights the conditions required to ensure the consistency of the prediction set under different schemes of interval censoring. Subsequently, we will extend our analysis to the estimation of oracle prediction sets $C^{*}\in\mathcal{C}$ consisting of multiple intervals. To measure the difference between two sets $A,B\subset\mathbb{R}$, we use the volume of their symmetric difference, denoted by $\mu(A\bigtriangleup B)$. The symmetric difference is defined as $A\bigtriangleup B=(A\setminus B)\cup(B\setminus A)$. This is commonly used in the literature on level set estimation and conformal inference. Another commonly used notion of set distance is the Hausdorff distance. In Appendix D, we discuss the relationship between these two metrics and why the Hausdorff distance is unsuitable in certain cases.
@@
where $K_{h}(x)=\frac{1}{h}K\left(\frac{x}{h}\right)$ for a choice of kernel smoothing function $K(\cdot)$ and bandwidth parameter $h$, see Tsybakov, (2009) and Hansen, (2008). In this case, $\mathcal{X}_{n}$ can be an expanding subset of $\mathcal{X}$ at a suitable rate to achieve uniform consistency as in Hansen, (2008), Theorem 8 (see discussion in Appendix B). Another candidate is the distributional regression estimator (see, e.g., Foresi and Peracchi,, 1995; Chernozhukov et al.,, 2013, 2021).
@@
Assumption 3.2 states that we cannot achieve the same coverage probability with a different interval $(t_{0}(x),t_{1}(x))$ that is not larger than $(\tau_{0}(x),\tau_{1}(x))$. Figure 1 shows a specific joint density of $Y^{L}$ and $Y^{U}$. The set $\{(a,b):t_{0}\leq a\leq b\leq t_{1}\}$ is a triangular region in $\mathbb{R}^{2}$. In Figure 1, the plotted triangular region corresponds to $t_{0}=-0.75$ and $t_{1}=0.5$. $P(t_{0},t_{1};x)$ will be the integral of the joint density over the triangular region. The first part of Assumption 3.2 thus requires that if either the triangular region is smaller or if we move the triangular region away from the optimal position, we will strictly lose coverage probability of $(Y^{L},Y^{U})$; and the second part is a regularity condition ensuring that the “coverage loss” is bounded away from 00 over the support of $X$. This rules out the case when $P(t_{0},t_{1};x)$ gets flatter around $(\tau_{0}(x),\tau_{1}(x))$ for some $x$.
Figure 1: Joint density of $Y^{L}$ and $Y^{U}$ when $Y\sim N(0,1)$, $\kappa^{L},\kappa^{U}\sim\text{Exp}(1)$. Integral of the density over the triangular region corresponds to $P(-0.75,0.5;x)$.
@@
Assumption 3.3 is a smoothness assumption on $P(t_{0},t_{1};x)$. In particular, we are assuming that by considering a slightly larger interval $[\tau_{0}-\varepsilon,\tau_{1}+\varepsilon]$ we can strictly increase our coverage probability. Assumption 3.3 also assumes an upper bound on the density, which can be violated if the conditional distribution of $(Y^{L},Y^{U})$ is discrete, which might happen under a fixed censoring scheme. This assumption is related to the $\gamma$-exponent condition in the level set literature, see Polonik, (1995) and Lei and Wasserman, (2014).
@@
Assumption 3.3 can be violated when there is fixed censoring, i.e., when there are some fixed brackets and we observe only the bracket an outcome falls in. Suppose the conditional distribution of $(Y^{L},Y^{U})$ can be decomposed into a continuous part and a discrete part, where the discrete part has support over the set of points $\{(L_{k},U_{k})\in\mathbb{R}^{2}:1\leq k\leq K\}$.
In the case of a mixture of random and fixed censoring, the regularity condition in Assumption 3.3 is violated. We impose the following assumption which slightly modifies the estimator to allow for non-smoothness of the conditional distribution.
@@
To see the reason why fixed censoring creates a problem and we should not solve the exact constraint, consider the following example.
@@
where $\bar{Y}^{U}=\frac{1}{n}\sum_{i}Y_{i}^{U}$. Since the lower bound for $(Y_{L},Y_{U})$ is 0 with probability one, it is easy to see that the lower bracket of $\check{C}$ is 00. Because
@@
the upper bracket of $\check{C}$ is $2$ if $P_{n}(0,1)<0.5$, i.e. there are less observations of $Y_{i}^{U}=1$ than $Y_{i}^{U}=2$ in the sample, which is equivalent to having the average $\bar{Y}^{U}>1.5$. Since $\bar{Y}^{U}$ has a scaled Binomial distribution $\bar{Y}^{U}\sim 1+n^{-1}\text{Binom}(n,0.5)$, and the median for a $\text{Binom}(n,p)$ distribution is between $\lfloor np\rfloor$ and $\lceil np\rceil$, we have that $P(\bar{Y}^{U}>1.5)\to 0.5$ as $n\to\infty$. This means that $P(\check{C}=[0,2])\to 0.5$ as $n\to\infty$ and $\check{C}$ is not consistent. On the other hand, for any $0.50<\psi_{n}<$ 0.5, $P_{n}(0,1)<0.5-\psi_{n}$ is equivalent to $\bar{Y}^{U}>1.5+\psi_{n}$, which has probability approaching 00 by the central limit theorem if $\sqrt{n}\psi_{n}\to\infty$ as the sample size $n\to\infty$. Solving the modified optimisation problem in (9) will then lead to a consistent estimator of the oracle prediction set in this example. Also note that the condition in (10) is a generalisation of the condition $\sqrt{n}\psi_{n}\to\infty$.
@@
    For any $\varepsilon>0$ and $P_{X}$-almost surely $x\in\mathcal{X}$, there exists $\delta=\delta(\varepsilon)>0$, such that for any set ${C}^{\prime}(x)\in\mathcal{C}$ with $\mu({C}^{\prime}(x))\leq\mu(C^{*}(x))$ and $\mu(C^{*}(x)\triangle C^{\prime}(x))>\varepsilon$, we have $P([Y^{L},Y^{U}]\subset{C}^{\prime}(x)\mid X=x)<1-\alpha-\delta$.
@@
    The oracle $C^{*}(x)\in\mathcal{C}_{\bar{M}}$, where $\mathcal{C}_{\bar{M}}=\{\sqcup_{m=1}^{M}[a_{m},b_{m}]:M^{\prime}\leq\bar{M}<\infty\}$, for some $\bar{M}<\infty$ and all $x\in\mathcal{X}$.
With the notation $P(C(x);x)=P([Y^{L},Y^{U}]\subset C(x)\mid X=x)$, and suppose that we have a uniform consistent estimator $P_{n}(C;x)$. We define the estimator $\hat{C}(x),x\in\mathcal{X}$, for the oracle prediction set $C^{*}$ allowing for potential fixed censoring.
@@
Appendix B states conditions under which a kernel smoothing estimator $P_{n}(C;x)$ is uniformly consistent and provides the corresponding convergence rate. This convergence rate can be used to inform the choice of $\psi_{n}$; in practice, it is often sufficient to simply set $\psi_{n}=0$. The uniformity is restricted to the class $\mathcal{C}_{M}$ rather than the larger class $\mathcal{C}$ because the proof relies on arguments that are closely tied to $\mathcal{C}_{M}$ being a VC class of sets, whereas $\mathcal{C}$ is not. Notably, knowledge of $\bar{M}$ is not required, as it suffices to select an $M$ that is sufficiently large. Even if $M$ is smaller than the true $\bar{M}$, the resulting prediction set can be viewed as an approximation for the oracle.
@@
4 Conformal inference with valued outcomes

### v4

@@ 2 Setup and definitions
A prediction set is defined as a set $C\subset\mathcal{X}\times\mathcal{Y}$, with the interpretation that, given a realisation of the predictor $X=x$, we predict that $Y$ is likely to fall inside the section $C(x)=\left\{y:(x,y)\in C\right\}$. In the absence of interval censoring, estimation of a prediction set that satisfies certain optimality criteria (i.e. the minimal prediction set with coverage guarantees) has been studied in the literature; see Lei and Wasserman (2014), Barber, Candes, Ramdas, and Tibshirani (2021) among others. As we discuss later in this section, the oracle prediction set without interval censoring is related to the upper level set of the density function of $Y$ given $X$ under point identification.
The situation is significantly different when we are faced with interval censoring and $P_{X,Y}$ is partially identified. In the rest of this section, we first define an optimality criterion for a prediction set under censoring, and analyse why the classical oracle prediction set (which corresponds to an upper level set) is no longer applicable under censoring. We then propose a feasible estimation procedure for the oracle prediction set. We discuss the potential reasons why the procedure can result in conservative prediction sets and show that the proposed procedure does not incur any unnecessary conservativeness.
There are two criteria for an oracle prediction set: validity and efficiency (see Vovk et al., 2009; Lei et al., 2013). Let $\mathcal{P}_{I}$ be the identified set of the joint distribution of $(X,Y)$ under censoring. That is, $\mathcal{P}_{I}$ is the set of joint distributions of $(X,Y)$ that are compatible with the joint distribution $P_{X,Y^{L},Y^{U}}$ and some censoring mechanism. A prediction set $C$ is valid if it satisfies the following coverage properties for a new sample $(X_{n+1},Y_{n+1})$ drawn from the same distribution $P$ at a given miscoverage level $\alpha$.
@@
When the model is point identified, $\mathcal{P}_{I}=\left\{P_{X,Y}\right\}$ is a singleton set, the conditions in Equations (2) and (3) reduce to the standard conditions of validity, see for example, Lei and Wasserman (2014). We will simply refer to these two conditions as marginal validity and conditional validity henceforth.
It is natural to require a prediction set to be valid, so that we have confidence it will contain the true value of the outcome with a certain probability. However, the requirement of validity by itself is vacuous. For example, the prediction set $C$ such that $C(x)=(-\infty,\infty)$, for all $x\in\mathcal{X}$, is technically valid at any miscoverage level, but it is also uninformative. Therefore, our target parameter of interest is an oracle prediction set, denoted by $C^{*}_{\mathcal{P}_{I}}$, which is defined as the minimal-volume prediction set that satisfies the validity conditions in Definition 2.1. For simplicity, we will omit the subscript and write $C^{*}=C^{*}_{\mathcal{P}_{I}}$ whenever the dependence on the identified set $\mathcal{P}_{I}$} is clear. Since the minimal-volume prediction set is intuitively the most informative among all valid prediction sets, this property is also called efficiency. [^1] This notion of efficiency, (see Vovk et al., 2009; Lei et al., 2013; Lei and Wasserman, 2014) differs from the usual definition, where an estimator is considered efficient if it has the smallest variance within a class of estimators. However, the concepts are related, as an estimator with a smaller variance generally leads to the construction of tighter and more informative confidence intervals.
@@
$\mathcal{C}=\left\{\bigsqcup_{m=1}^{M}[a_{m},b_{m}]:a_{m}<b_{m}<a_{m+1},M<\infty\right\}$
denotes the collection of all finite disjoint unions of closed intervals, and the symbol $\sqcup$ indicates that the union is disjoint. The following oracle prediction set under partial identification will be our estimation target.
@@
The oracle prediction set under point identification is $C^{*}_{P_{Y|X}}=\left\{(x,y):x\in\mathcal{X},y\in L(x,\lambda_{x}^{\alpha})\right\}$, which is both conditionally and marginally valid (see Lei and Wasserman (2014)). For a given $x$, it collects the values of $y$ with sufficiently high density and intuitively assembles the set of $y$’s with the highest probability for a given volume. Once this is calibrated to yield the desired miscoverage level, one obtains a minimal-volume set. Estimating the oracle prediction set $C^{*}_{P_{Y|X}}$ therefore reduces to the problem of estimating the upper level sets $L(x,\lambda)$ of the conditional density $p(y\mid x)$ and selecting an appropriate threshold parameter.
However, this characterisation of the oracle prediction set in terms of level sets is no longer applicable under interval censoring, as $P_{Y\mid X}$ is not identified. In fact, the next proposition states that no meaningful bounds can be obtained for the level set of the conditional density of $Y$ given $X$ under interval censoring. For brevity, we temporarily drop the conditioning predictors $X$ in the following proposition. (It can be interpreted as conditional on $X=x$ for $x\in\mathcal{X}$.)
@@
Lemma 2.2 (Theorem SIR-2.3 in Molinari (2020)).
@@
for a disjoint union of intervals $C\in\mathcal{C}$. As a result, a prediction set that satisfies the feasible version of the constraint will have a coverage guarantee since
@@
The above lemma shows that even though the oracle prediction set can be conservative, i.e. overcovering $Y$ for the distribution actually generating the data, the conservativeness is only due to partial identification and the conditional distribution of $(Y^{L},Y^{U})$ and not because the constraint we focus on differs from the original one.
@@
The oracle prediction set $C^{*}$ is the solution to the following feasible optimisation problem for $x$ in the support of $X$,
@@
A natural estimator $\hat{C}$ for the oracle prediction set can be obtained by solving the sample version of the optimisation problem above. The next two sections study the consistency of such estimators for the oracle prediction set and provide a modification of the estimated prediction set that has finite sample coverage guarantees using conformal inference.
@@ 3 Consistency
We first consider the estimation of the oracle prediction interval $C^{*}_{I}=[\tau_{0}(x),\tau_{1}(x)]$, for $x\in\mathcal{X}$. This analysis highlights the conditions required to ensure the consistency of the prediction set under different schemes of interval censoring. Subsequently, we will extend our analysis to the estimation of oracle prediction sets $C^{*}\in\mathcal{C}$ consisting of multiple intervals. To measure the difference between two sets $A,B\subset\mathbb{R}$, we use the volume of their symmetric difference, denoted by $\mu(A\bigtriangleup B)$. The symmetric difference is defined as $A\bigtriangleup B=(A\setminus B)\cup(B\setminus A)$. This is commonly used in the literature on level set estimation and conformal inference. Another commonly used notion of set distance is the Hausdorff distance. In Appendix D, we discuss the relationship between these two metrics and why the Hausdorff distance is unsuitable in certain pathological cases.
@@
where $K_{h}(x)=\frac{1}{h}K\left(\frac{x}{h}\right)$ for a choice of kernel smoothing function $K(\cdot)$ and bandwidth parameter $h$, see Tsybakov (2009) and Hansen (2008). In this case, $\mathcal{X}_{n}$ can be an expanding subset of $\mathcal{X}$ at a suitable rate to achieve uniform consistency as in Hansen (2008), Theorem 8 (see discussion in Appendix B). Another candidate is the distributional regression estimator (see, e.g., Foresi and Peracchi, 1995; Chernozhukov et al., 2013, 2021).
@@
Assumption 3.2 states that we cannot achieve the same coverage probability with a different interval $(t_{0}(x),t_{1}(x))$ that is not larger than $(\tau_{0}(x),\tau_{1}(x))$. Figure 1 shows a specific joint density of $Y^{L}$ and $Y^{U}$. The set $\{(a,b):t_{0}\leq a\leq b\leq t_{1}\}$ is a triangular region in $\mathbb{R}^{2}$. In Figure 1, the plotted triangular region corresponds to $t_{0}=-0.75$ and $t_{1}=0.5$. $P(t_{0},t_{1};x)$ will be the integral of the joint density over the triangular region. The first part of Assumption 3.2 thus requires that if either the triangular region is smaller or if we move the triangular region away from the optimal position, we will strictly lose coverage probability of $(Y^{L},Y^{U})$; and the second part is a regularity condition ensuring that the “coverage loss” is bounded away from 0 over the support of $X$. This rules out the case when $P(t_{0},t_{1};x)$ gets flatter around $(\tau_{0}(x),\tau_{1}(x))$ for some $x$.
Figure 1: Joint density of $Y^{L}$ and $Y^{U}$ when $Y\sim N(0,1)$, $\kappa^{L},\kappa^{U}\sim\text{Exp}(1)$. Integral of the density over the triangular region corresponds to P(0.75,0.5;x)P(-0.75,0.5;x).
@@
Assumption 3.3 is a smoothness assumption on $P(t_{0},t_{1};x)$. In particular, we are assuming that by considering a slightly larger interval $\tau_{0}-\varepsilon,\tau_{1}+\varepsilon]$ we can strictly increase our coverage probability. Assumption 3.3 also assumes an upper bound on the density, which can be violated if the conditional distribution of $(Y^{L},Y^{U})$ is discrete, which might happen under a fixed censoring scheme. This assumption is related to the $\gamma$-exponent condition in the level set literature, see Polonik (1995) and Lei and Wasserman (2014).
@@
Assumption 3.3 can be violated when there is fixed censoring, i.e., when there are some fixed brackets, and we observe only the bracket in which an outcome falls. Suppose the conditional distribution of $(Y^{L},Y^{U})$ can be decomposed into a continuous part and a discrete part, where the discrete part has support over the set of points $\{(L_{k},U_{k})\in\mathbb{R}^{2}:1\leq k\leq K\}$.
In the case of a mixture of random and fixed censoring, the regularity condition in Assumption 3.3 is violated. We impose the following assumption, which slightly modifies the estimator to allow for non-smoothness of the conditional distribution.
@@
To see the reason why fixed censoring creates a problem, and we should not solve the exact constraint, consider the following example.
@@
where $\bar{Y}^{U}=\frac{1}{n}\sum_{i}Y_{i}^{U}$. Since the lower bound for $(Y_{L},Y_{U})$ is 0 with probability one, it is easy to see that the lower bracket of $\check{C}$ is 0. Because
@@
the upper bracket of $\check{C}$ is $2$ if $P_{n}(0,1)<0.5$, i.e. there are fewer observations of $Y_{i}^{U}=1$ than $Y_{i}^{U}=2$ in the sample, which is equivalent to having the average $\bar{Y}^{U}>1.5$. Since $\bar{Y}^{U}$ has a scaled Binomial distribution $\bar{Y}^{U}\sim 1+n^{-1}\text{Binom}(n,0.5)$, and the median for a $\text{Binom}(n,p)$ distribution is between $\lfloor np\rfloor$ and $\lceil np\rceil$, we have that $P(\bar{Y}^{U}>1.5)\to 0.5$ as $n\to\infty$. This means that $P(\check{C}=[0,2])\to 0.5$ as $n\to\infty$ and $\check{C}$ is not consistent. On the other hand, for any $0<\psi_{n}<0.5$, $P_{n}(0,1)<0.5-\psi_{n}$ is equivalent to $\bar{Y}^{U}>1.5+\psi_{n}$, which has probability approaching 0 by the central limit theorem if $\sqrt{n}\psi_{n}\to\infty$ as the sample size $n\to\infty$. Solving the modified optimisation problem in (9) will then lead to a consistent estimator of the oracle prediction set in this example. Also note that the condition in (10) is a generalisation of the condition $\sqrt{n}\psi_{n}\to\infty$.
@@
    There exists a unique solution $C^{*}$ to 2.4. For any $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$, such that for $P_{X}$-almost surely $x\in\mathcal{X}$ and any set ${C}^{\prime}(x)\in\mathcal{C}$ with $\mu({C}^{\prime}(x))\leq\mu(C^{*}(x))$ and $\mu(C^{*}(x)\triangle C^{\prime}(x))>\varepsilon$, we have $P([Y^{L},Y^{U}]\subset{C}^{\prime}(x)\mid X=x)<1-\alpha-\delta$.
@@
    The oracle $C^{*}(x)\in\mathcal{C}_{\bar{M}}$, where $\mathcal{C}_{\bar{M}}=\{\sqcup_{m=1}^{M}[a_{m},b_{m}]:a_{m}<b_{m}<a_{m+1},\,M\leq\bar{M}<\infty\}$, for some $\bar{M}<\infty$ and almost surely $x\in\mathcal{X}$.
With the notation $P(C(x);x)=P([Y^{L},Y^{U}]\subset C(x)\mid X=x)$, and suppose that we have a uniform consistent estimator $P_{n}(C;x)$. We define the estimator $\hat{C}(x),x\in\mathcal{X}$, for the oracle prediction set $C^{*}$, allowing for potential fixed censoring.
@@
Appendix B states conditions under which a kernel smoothing estimator $P_{n}(C;x)$ is uniformly consistent and provides the corresponding convergence rate. This convergence rate can be used to inform the choice of $\psi_{n}$. Empirically, varying $\psi_{n}$ between 0 and small positive values produces negligible changes in the results. The uniformity is restricted to the class $\mathcal{C}_{M}$ rather than the larger class $\mathcal{C}$ because the proof relies on arguments that are closely tied to $\mathcal{C}_{M}$ being a VC class of sets. Notably, knowledge of $\bar{M}$ is not required, as it suffices to select an $M$ that is sufficiently large. Even if $M$ is smaller than the true $\bar{M}$, the resulting prediction set can be viewed as an approximation for the oracle prediction set.
@@
4 Conformal inference with interval outcomes